40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4273     Entry time: Fri Feb 11 09:27:03 2011
Author: steve 
Type: Configuration 
Category: VAC 
Subject: RGA scan 

The RGA scan is normal at day 52 of this pump down.

Light power BS 1064nm ~25mW, ETMX 532nm ~5mW

Attachment 1: rgascan20110211.jpg  124 kB  | Hide | Hide all | Show all
rgascan20110211.jpg
Attachment 2: pressurepd70.jpg  81 kB  | Show | Hide all | Show all
ELOG V3.1.3-